a+b+c=11ab+BC+ca=12Then what is the value of ab(a+b)+bc(b+c)+ac(a+c)

Grade:12th pass

1 Answers

sahil
6 years ago
see a+b+c=11 and ab+bc+ca=12 then ab(a+b)+bc(b+c)+ca(c+a)=ab(11-c)+bc(11-a)+ca(11-b)=11(ab+bc+ca)-3abc=132-3abc if this is answer let me know by clicking on approval option and if -3abc value can be calculated please tell so i could try to find -3abc